OpenStudy (anonymous):

What is the inverse of 70x^(3/4) Please explain each step.

OpenStudy (aum):

\[ y = 70x^{3/4} \\ \text{Solve for x:} \\ \text{Divide by 70:} \\ \frac{y}{70} = x^{3/4} \\ \text{Raise both sides to 4/3:} \\ ?? \]

OpenStudy (anonymous):

Does it matter if you switch the x and why first withing the function, because I noticed that you didn't.

OpenStudy (aum):

You can do it at the beginning or at the end. You will get the same result.

OpenStudy (anonymous):

why did you raise it to the 4/3 power and how does it cancel out the ^3/4

OpenStudy (anonymous):

Yeah it doesn't really matter, but I like to do this... To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

Cancels the right side exponent

OpenStudy (aum):

\[ (x^{3/4})^{4/3} = x^1 = x \]

OpenStudy (anonymous):

Go aum! :D

OpenStudy (anonymous):

Ok so now i understand that, would the answer be---- (x^4/3)/70^4/3

OpenStudy (aum):

\[ y = 70x^{3/4} \\ \text{Switch x and y:} \\ x = 70y^{3/4} \\ \text{Solve for y:} \\ \text{Divide by 70:} \\ \frac{x}{70} = y^{3/4} \\ \text{Raise both sides to 4/3:} \\ \left(\frac{x}{70}\right)^{4/3} = (y^{3/4})^{4/3} = y^{3/4*4/3} = y^1 = y \\ y = \left(\frac{x}{70}\right)^{4/3} \\ f^{-1}(x) = \left(\frac{x}{70}\right)^{4/3} \]

OpenStudy (anonymous):

OK THANKS SO MUCH AUM!!! I understand. :D

OpenStudy (aum):

You are welcome.